# The Taylor-Wiles method for coherent cohomology, II

### v2

@@ 2.3. Ramification and types
The modules $H^{q}_{!}(R\Gamma({\mathbb{S}},{\mathcal{F}}_{K}))$ of coherent cohomology give rise to Galois representations that may be ramified at places $v$ where $K_{v}$ is not hyperspecial maximal. In order to restrict attention to minimal deformation problems, we introduce spaces of cohomology with coefficients in semisimple Bushnell-Kutzko types [BK99]. This was explained in the unpublished manuscript [30], in the case of supercuspidal types, but seems have never been used subsequently, although the definition we give below of automorphic forms with coefficients in Bushnell-Kutzko types is essentially the one that appeared in [30] as well as [63, 37]; for automorphic forms on $GL(2)$ it has been standard for some time. A version based on finite-dimensional representations of the multiplicative groups of local division algebras is developed in [10, p. 97]; this is essentially equivalent for purposes of proving modularity but it cannot be applied to prove that the coherent cohomology is free over the localized Hecke algebra, which is our main theorem.
@@ 2.4. Minimality condition
For applications to the Taylor-Wiles method, we add the following hypotheses on the Galois representations attached to the type data. For each $v$, we choose a $\pi$ in the inertial equivalence class corresponding to the quadruple $(K_{v}^{+},K_{v},\mbox{ю}_{v},\Lambda_{\mbox{ю}_{v}})$ – in other words, ${\mathrm{Hom}}_{K_{v}^{+}}(\mbox{ю}_{v},\pi)\neq 0$. Let ${\mathcal{L}}_{v}(\pi)$ denote the $n$-dimensional representation of the Weil-Deligne group $W_{v}$ of $F^{+}_{v}$ attached to $\pi$ by the local Langlands correspondence, with coefficients in the $p$-adic integer ring ${\mathcal{O}}$, and let ¯$\overline{{\mathcal{L}}}_{v}(\pi)$ denote its reduction modulo the maximal ideal ${\mathfrak{m}}_{\mathcal{O}}$. We attach to ${\mathcal{L}}_{v}(\pi)$ a deformation problem $\mathcal{D}_{v}$ over ${\mathcal{O}}$, in the sense of [10, Definition 2.2.2], by the condition that, ${\mathcal{L}}_{v}(\pi^{\prime})$ is a deformation of ${\mathcal{L}}_{v}(\pi)$ of type $\mathcal{D}_{v}$ if and only if ${\mathrm{Hom}}_{K_{v}^{+}}(\mbox{ю}_{v},\pi^{\prime})\neq 0$. To $\mathcal{D}_{v}$ is attached a subspace $L_{v}\subset H^{1}(W_{v},{\mathrm{ad}}(\overline{{\mathcal{L}}}_{v}(\pi)))$, as in [10, §2.2], and we assume that the problem $\mathcal{D}_{v}$ is minimal in the sense that
@@ 3.1. Review of the theorem of Nakajima
We begin with a simple generalization of a special case of the theorem of Nakajima [49, Theorem 2] used in [27].
@@
The following corollary is then proved just as in [27, Corollary 3.5]:
@@
The following Corollary extends [27, Proposition 3.7] to the noncompact case.
@@
$H^{i_{0}(w,\mu)}(S_{\Delta}(Q),{\mathcal{E}}_{w}(\mu)^{\rm can})_{{\mathfrak{m}}}$
@@
$H^{i}(\mathbb{S}_{\Delta}(Q),{\mathcal{E}}_{w}(\mu)^{\rm can})_{{\mathfrak{m}}}=0\quad\text{ if }i\neq i_{0}(w,\mu).$
@@
In view of Theorem 15, we can apply the same proof as in [27, Proposition 3.7]. ∎
@@ 4. Application of the Taylor-Wiles method
Note that by Proposition  27, these are torsion-free ${\mathcal{O}}$-modules. The Hecke algebras ${\mathbb{T}}_{\emptyset,w},{\mathbb{T}}_{0,Q,w},{\mathbb{T}}_{\Delta,Q,w}$ are defined as in [27, p. 137] with respect to these modules as the images of the relevant Hecke algebras ${\mathbb{T}}_{S,K}(w,\mu)$, initially defined as endomorphisms of the perfect complex, on the cohomology modules. Given a homomorphism
@@
We also assume the residual representation $\bar{\rho}_{\nu}$ to be absolutely irreducible444needed for the construction of the sets $Q_{M}$ in Prop. 4.4 by Thorne. and adequate in the sense of [60]. Let ${\mathfrak{m}}={\mathfrak{m}}(\bar{\rho}_{\nu})\subset{\mathbb{T}}_{\emptyset,w}(\mu)$ be the corresponding maximal ideal, and let $H_{\bar{\rho}_{\nu},\emptyset}$ and ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$ denote the localizations at ${\mathfrak{m}}$. Suppose that the deformation conditions are minimally ramified at all primes $v\in S$ not dividing $p$. In particular, for $v\in S$, we assume the deformation condition in ${\mathcal{S}}$ at $v$ is unrestricted and minimal (cf. [27, §6.9]).
@@
The reader can check that, since $p$ is odd, the condition on $v\in S(F^{+})$ is not strictly necessary; this is the point of [10, Lemma 2.3.3], whose application is buried at the end of the proof of Lemma 2.3.4 of the same paper. In order to avoid distraction we add this as an extra condition.
@@
The Galois Hypothesis 4.3 follows as in the proofs of Corollaries 5.3 and 5.4 in [27] in light of our Theorem 24 (ii). Note that ${\mathbb{T}}_{\emptyset,w}(\mu)$ may not be reduced; for that reason in all proofs we use the reduced localizations. Hypothesis 4.4.8 is verified in Lemma 6.1 therein. It remains to check the correct analog of Hypothesis 4.4.9 for adequate residual representations.
@@
Finally, the modifications necessary for incorporating the type data are the same as in the proof of [10, Theorem 3.5.1], which specifically comes down to the use of the minimality hypothesis in [10, Proposition 2.5.9]. ∎
@@ 5.1. Topological and de Rham cohomology
Theorem 36 asserts that the coherent cohomology module $H^{i_{0}}({\mathbb{S}}_{K,\Sigma},{\mathcal{E}}_{w}(\mu)^{\rm can})$ localized at the maximal ideal ${\mathfrak{m}}={\mathfrak{m}}(\bar{\rho}_{\nu})$ is free over the localized Hecke algebra ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$, provided $p$ and the weight $\mu$ satisfy the stated inequalities. As in [27, §7], this implies that analogous localized modules in $p$-adic étale and de Rham cohomology are also free over ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$. The proofs are identical, so we merely state the results.
Let now $W=W_{\mu}$ be the finite-dimensional irreducible representation of $G$ with highest weight $\mu$. As in [27, (7.1.2)] we always assume $\mu$ to be $p$-small. Let the $p$-adic place $v$, the $p$-adic integer ring ${\mathcal{O}}$, the hyperspecial maximal compact subgroup $K_{p}\subset G({{\mathbb{Q}}_{p}})$, and the $K_{p}$-invariant lattice $W_{\mu}({\mathcal{O}})\subset W(Frac({\mathcal{O}}))$ be as in [27, §7.1]. Let $\tilde{W}_{\mu,B}(\mathcal{O})$ denote the topological local system in ${\mathcal{O}}$-modules on $S_{K}(G,X)(\mathbb{C})$, as in [27, §7.3]. Topological cohomology is computed by the complex
@@
The de Rham version ${\mathbb{T}}_{\emptyset,dR}(\mu)$ of ${\mathbb{T}}_{\emptyset,B}(\mu)$ requires a bit more work to define. Let ${\mathcal{E}}_{G,{\mathcal{O}}}$ be the exact monoidal functor defined in [47, Lemma 1.20] from the tensor category of representations of the group scheme $G_{{\mathbb{Z}}_{p}}$ over ${\mathcal{O}}$ to the category of locally free coherent sheaves on ${\mathbb{S}}_{K}$ with integrable connection and compatible $G({\mathbf{A}}_{f}^{p})$ action covering the natural action on the family ${\mathbb{S}}_{K\times K^{p}}$. Let $W_{\mu,{\mathcal{O}}}$ denote the representation of $G_{{\mathbb{Z}}_{p}}$ introduced in [27, §7.1], so that the ${\mathcal{O}}$-lattice $W_{\mu}({\mathcal{O}})$ is just $W_{\mu,{\mathcal{O}}}({\mathcal{O}})$. We let
@@
see [48, §4.3], or [48, §7.4] for a reformulation in terms of dual BGG complexes. We define ${\mathbb{T}}_{\emptyset,dR}(\mu)\subset End\left(R\Gamma_{dR,log}({\mathbb{S}}_{K},W_{\mathcal{O}}))\right)$ again to be the ${\mathcal{O}}$-subalgebra generated by the algebraic Hecke correspondences $T(t)$.
@@
is a free ${\mathcal{O}}$-module of finite rank. Moreover, there is a natural decreasing (Hodge) filtration $F^{\bullet}H^{d_{V}}_{dR,log}({\mathbb{S}}_{K},W_{\mathcal{O}}))_{\bar{\rho}_{\nu}}$ by ${\mathcal{O}}$-direct summands satisfying the analogue of [27, Theorem 7.2.2].
(ii) Assume in addition the inequality $|\mu_{comp}|\leq p-2$, as in the statement of [27, Theorem 7.3.3]. Then
@@
The next two theorems are then proved exactly as in [27, §7.2, §7.3].
@@
Assume Hypotheses 39 and the inequality $|\mu_{comp}|\leq p-2$, as in the statement of [27, Theorem 7.3.3]. Then there is an isomorphism
@@ 5.3. Exterior powers of modules over Hecke algebras
The congruence ideal $C(\nu,{\mathbb{T}}_{V})=C(\pi_{\nu},{\mathbb{T}}_{V})\subset{\mathbb{T}}_{V}$ is defined as in [14, Definition 6.7.2]. More precisely, it is the annihilator of
@@ 6. Ordinary modular forms
The article [14] constructs $p$-adic $L$-functions as elements of Hida’s ordinary Hecke algebra when the latter, localized at an appropriate maximal ideal, is known to satisfy a Gorenstein hypothesis [14, §6.7.8, §7.3.2]. Here we show how to derive this hypothesis from Theorem 36, when $p$ is sufficiently large and the maximal ideal is as in the statement of that theorem. The first results of this type are due to Hida [34] and Tilouine [62]; our method is essentially the same as Tilouine’s.
We make use of the notation of [14] without comment. Thus $\pi$ is an anti-holomorphic cuspidal automorphic representation of the unitary similitude group $G$ (denoted $G_{1}$ in loc. cit.) and ${\mathbb{T}}={\mathbb{T}}_{\pi}$ is the localization of Hida’s ordinary Hecke algebra at the corresponding maximal ideal ${\mathfrak{m}}_{\pi}$. This ideal determines a connected component of weight space, which includes the weight $\kappa(\pi)$ of the holomorphic modular form corresponding to $\pi$ (more conventionally, to the contragredient of $\pi$, a holomorphic cuspidal automorphic representation), which we are allowed to vary. Theorem 36 can be applied provided $\pi$ can be chosen of a weight that satisfies the inequalities in Proposition 27 as well as those of Hida’s Control Theorem, cited as [14, Theorem 7.2.1]. We make these conditions more precise in the following Theorem. We say two weights $\mu,\mu^{\prime}$ are congruent modulo $p$ if they define the same characters on the (finite group of) points of the special fiber of the torus $T$ over the residue field $k$ of ${\mathcal{O}}$.
@@
The notation $\bar{\rho}_{\pi}$ corresponds to the notation $\bar{\rho}_{\nu}$ in the statement of Theorem 36. Applying that theorem, the hypotheses imply that the localized module $S_{\kappa(\pi)}(K)_{\pi}$ is free over the localized Hecke algebra, which is a local complete intersection. We claim that it suffices to show that Hida’s Control Theorem, as stated in [35, Theorem 7.1], applies to $S_{\kappa(\pi)}(K)_{\pi}$. Indeed, the Hecke algebra ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$ in weight $\kappa(\pi)$ is the quotient of ${\mathbb{T}}_{\pi}$ by the regular sequence corresponding to the weight $\kappa(\pi)$. Moreover, we are concerned with coherent cohomology in degree 00, so by Koecher’s principle the reduced localization hypothesis in Theorem 36 is superfluous.
Now it remains to verify that Hida’s Control Theorem does apply to weight $\kappa(\pi)$ when $\mu^{\prime}$ satisfies the regularity condition (ii) of Proposition 27. In fact, Boxer and Pilloni have proved a classicality theorem for overconvergent modular forms of small slope [7, Theorem 1.0.15 (3)] that asserts that overconvergent modular forms of weight $\mu$ on Shimura varieties of abelian type are classical if they satisfy a small slope condition and if $\mu^{\prime}$ satisfies the regularity condition (ii) of Proposition 27. Since ordinary modular forms automatically satisfy the small slope condition (see [7, Remark 1.0.6]), this completes the proof.

### v3

@@ 2.3. Ramification and types
The modules $H^{q}_{!}(R\Gamma({\mathbb{S}},{\mathcal{F}}_{K}))$ of coherent cohomology give rise to Galois representations that may be ramified at places $v$ where $K_{v}$ is not hyperspecial maximal. In order to restrict attention to minimal deformation problems, we introduce spaces of cohomology with coefficients in semisimple Bushnell-Kutzko types [BK99]. This was explained in the unpublished manuscript [31], in the case of supercuspidal types, but seems have never been used subsequently, although the definition we give below of automorphic forms with coefficients in Bushnell-Kutzko types is essentially the one that appeared in [31] as well as [65, 38]; for automorphic forms on $GL(2)$ it has been standard for some time. A version based on finite-dimensional representations of the multiplicative groups of local division algebras is developed in [11, p. 97]; this is essentially equivalent for purposes of proving modularity but it cannot be applied to prove that the coherent cohomology is free over the localized Hecke algebra, which is our main theorem.
@@ 2.4. Minimality condition
For applications to the Taylor-Wiles method, we add the following hypotheses on the Galois representations attached to the type data. For each $v$, we choose a $\pi$ in the inertial equivalence class corresponding to the quadruple $(K_{v}^{+},K_{v},\mbox{ю}_{v},\Lambda_{\mbox{ю}_{v}})$ – in other words, ${\mathrm{Hom}}_{K_{v}^{+}}(\mbox{ю}_{v},\pi)\neq 0$. Let ${\mathcal{L}}_{v}(\pi)$ denote the $n$-dimensional representation of the Weil-Deligne group $W_{v}$ of $F^{+}_{v}$ attached to $\pi$ by the local Langlands correspondence, with coefficients in the $p$-adic integer ring ${\mathcal{O}}$, and let ¯$\overline{{\mathcal{L}}}_{v}(\pi)$ denote its reduction modulo the maximal ideal ${\mathfrak{m}}_{\mathcal{O}}$. We attach to ${\mathcal{L}}_{v}(\pi)$ a deformation problem $\mathcal{D}_{v}$ over ${\mathcal{O}}$, in the sense of [11, Definition 2.2.2], by the condition that, ${\mathcal{L}}_{v}(\pi^{\prime})$ is a deformation of ${\mathcal{L}}_{v}(\pi)$ of type $\mathcal{D}_{v}$ if and only if ${\mathrm{Hom}}_{K_{v}^{+}}(\mbox{ю}_{v},\pi^{\prime})\neq 0$. To $\mathcal{D}_{v}$ is attached a subspace $L_{v}\subset H^{1}(W_{v},{\mathrm{ad}}(\overline{{\mathcal{L}}}_{v}(\pi)))$, as in [11, §2.2], and we assume that the problem $\mathcal{D}_{v}$ is minimal in the sense that
@@ 3.1. Review of the theorem of Nakajima
We begin with a simple generalization of a special case of the theorem of Nakajima [51, Theorem 2] used in [28].
@@
The following corollary is then proved just as in [28, Corollary 3.5]:
@@
The following Corollary extends [28, Proposition 3.7] to the noncompact case.
@@
$H^{i_{0}(w,\mu)}(\mathbb{S}_{\Delta,Q,\Sigma},{\mathcal{E}}_{w}(\mu)^{\rm can})_{{\mathfrak{m}}}$
@@
$H^{i}(\mathbb{S}_{\Delta,Q,\Sigma},{\mathcal{E}}_{w}(\mu)^{\rm can})_{{\mathfrak{m}}}=0\quad\text{ if }i\neq i_{0}(w,\mu).$
@@
In view of Theorem 15, we can apply the same proof as in [28, Proposition 3.7]. ∎
@@ 4. Application of the Taylor-Wiles method
Note that by Proposition  27, these are torsion-free ${\mathcal{O}}$-modules. The Hecke algebras ${\mathbb{T}}_{\emptyset,w},{\mathbb{T}}_{0,Q,w},{\mathbb{T}}_{\Delta,Q,w}$ are defined as in [28, p. 137] with respect to these modules as the images of the relevant Hecke algebras ${\mathbb{T}}_{S,K}(w,\mu)$, initially defined as endomorphisms of the perfect complex, on the cohomology modules. Given a homomorphism
@@
We also assume the residual representation $\bar{\rho}_{\nu}$ to be absolutely irreducible444needed for the construction of the sets $Q_{M}$ in Prop. 4.4 by Thorne. and adequate in the sense of [62]. Let ${\mathfrak{m}}={\mathfrak{m}}(\bar{\rho}_{\nu})\subset{\mathbb{T}}_{\emptyset,w}(\mu)$ be the corresponding maximal ideal, and let $H_{\bar{\rho}_{\nu},\emptyset}$ and ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$ denote the localizations at ${\mathfrak{m}}$. Suppose that the deformation conditions are minimally ramified at all primes $v\in S$ not dividing $p$. In particular, for $v\in S$, we assume the deformation condition in ${\mathcal{S}}$ at $v$ is unrestricted and minimal (cf. [28, §6.9]).
@@
The reader can check that, since $p$ is odd, the condition on $v\in S(F^{+})$ is not strictly necessary; this is the point of [11, Lemma 2.3.3], whose application is buried at the end of the proof of Lemma 2.3.4 of the same paper. In order to avoid distraction we add this as an extra condition.
@@
The Galois Hypothesis 4.3 follows as in the proofs of Corollaries 5.3 and 5.4 in [28] in light of our Theorem 24 (ii). Hypothesis 4.4.8 is verified in Lemma 6.1 therein. It remains to check the correct analog of Hypothesis 4.4.9 for adequate residual representations.
@@
Finally, the modifications necessary for incorporating the type data are the same as in the proof of [11, Theorem 3.5.1], which specifically comes down to the use of the minimality hypothesis in [11, Proposition 2.5.9]. ∎
@@ 5.1. Topological and de Rham cohomology
Theorem 36 asserts that the coherent cohomology module $H^{i_{0}}({\mathbb{S}}_{K,\Sigma},{\mathcal{E}}_{w}(\mu)^{\rm can})$ localized at the maximal ideal ${\mathfrak{m}}={\mathfrak{m}}(\bar{\rho}_{\nu})$ is free over the localized Hecke algebra ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$, provided $p$ and the weight $\mu$ satisfy the stated inequalities. As in [28, §7], this implies that analogous localized modules in $p$-adic étale and de Rham cohomology are also free over ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$. The proofs are identical, so we merely state the results.
Let now $W=W_{\mu}$ be the finite-dimensional irreducible representation of $G$ with highest weight $\mu$. As in [28, (7.1.2)] we always assume $\mu$ to be $p$-small. Let the $p$-adic place $v$, the $p$-adic integer ring ${\mathcal{O}}$, the hyperspecial maximal compact subgroup $K_{p}\subset G({{\mathbb{Q}}_{p}})$, and the $K_{p}$-invariant lattice $W_{\mu}({\mathcal{O}})\subset W(Frac({\mathcal{O}}))$ be as in [28, §7.1]. Let $\tilde{W}_{\mu,B}(\mathcal{O})$ denote the topological local system in ${\mathcal{O}}$-modules on $S_{K}(G,X)(\mathbb{C})$, as in [28, §7.3]. Topological cohomology is computed by the complex
@@
The de Rham version ${\mathbb{T}}_{\emptyset,dR}(\mu)$ of ${\mathbb{T}}_{\emptyset,B}(\mu)$ requires a bit more work to define. Let ${\mathcal{E}}_{G,{\mathcal{O}}}$ be the exact monoidal functor defined in [49, Lemma 1.20] from the tensor category of representations of the group scheme $G_{{\mathbb{Z}}_{p}}$ over ${\mathcal{O}}$ to the category of locally free coherent sheaves on ${\mathbb{S}}_{K}$ with integrable connection and compatible $G({\mathbf{A}}_{f}^{p})$ action covering the natural action on the family ${\mathbb{S}}_{K\times K^{p}}$. Let $W_{\mu,{\mathcal{O}}}$ denote the representation of $G_{{\mathbb{Z}}_{p}}$ introduced in [28, §7.1], so that the ${\mathcal{O}}$-lattice $W_{\mu}({\mathcal{O}})$ is just $W_{\mu,{\mathcal{O}}}({\mathcal{O}})$. We let
@@
see [50, §4.3], or [50, §7.4] for a reformulation in terms of dual BGG complexes. We define ${\mathbb{T}}_{\emptyset,dR}(\mu)\subset End\left(R\Gamma_{dR,log}({\mathbb{S}}_{K},W_{\mathcal{O}}))\right)$ again to be the ${\mathcal{O}}$-subalgebra generated by the algebraic Hecke correspondences $T(t)$.
@@
is a free ${\mathcal{O}}$-module of finite rank. Moreover, there is a natural decreasing (Hodge) filtration $F^{\bullet}H^{d_{V}}_{dR,log}({\mathbb{S}}_{K},W_{\mathcal{O}}))_{\bar{\rho}_{\nu}}$ by ${\mathcal{O}}$-direct summands satisfying the analogue of [28, Theorem 7.2.2].
(ii) Assume in addition the inequality $|\mu_{comp}|\leq p-2$, as in the statement of [28, Theorem 7.3.3]. Then
@@
The next two theorems are then proved exactly as in [28, §7.2, §7.3].
@@
Assume Hypotheses 39 and the inequality $|\mu_{comp}|\leq p-2$, as in the statement of [28, Theorem 7.3.3]. Then there is an isomorphism
@@ 5.3. Exterior powers of modules over Hecke algebras
The congruence ideal $C(\nu,{\mathbb{T}}_{V})=C(\pi_{\nu},{\mathbb{T}}_{V})\subset{\mathbb{T}}_{V}$ is defined as in [15, Definition 6.7.2]. More precisely, it is the annihilator of
@@ 6. Ordinary modular forms
The article [15] constructs $p$-adic $L$-functions as elements of Hida’s ordinary Hecke algebra when the latter, localized at an appropriate maximal ideal, is known to satisfy a Gorenstein hypothesis [15, §6.7.8, §7.3.2]. Here we show how to derive this hypothesis from Theorem 36, when $p$ is sufficiently large and the maximal ideal is as in the statement of that theorem. The first results of this type are due to Hida [35] and Tilouine [64]; our method is essentially the same as Tilouine’s.
We make use of the notation of [15] without comment. Thus $\pi$ is an anti-holomorphic cuspidal automorphic representation of the unitary similitude group $G$ (denoted $G_{1}$ in loc. cit.) and ${\mathbb{T}}={\mathbb{T}}_{\pi}$ is the localization of Hida’s ordinary Hecke algebra at the corresponding maximal ideal ${\mathfrak{m}}_{\pi}$. This ideal determines a connected component of weight space, which includes the weight $\kappa(\pi)$ of the holomorphic modular form corresponding to $\pi$ (more conventionally, to the contragredient of $\pi$, a holomorphic cuspidal automorphic representation), which we are allowed to vary. Theorem 36 can be applied provided $\pi$ can be chosen of a weight that satisfies the inequalities in Proposition 27 as well as those of Hida’s Control Theorem, cited as [15, Theorem 7.2.1]. We make these conditions more precise in the following Theorem. We say two weights $\mu,\mu^{\prime}$ are congruent modulo $p$ if they define the same characters on the (finite group of) points of the special fiber of the torus $T$ over the residue field $k$ of ${\mathcal{O}}$.
@@
The notation $\bar{\rho}_{\pi}$ corresponds to the notation $\bar{\rho}_{\nu}$ in the statement of Theorem 36. Applying that theorem, the hypotheses imply that the localized module $S_{\kappa(\pi)}(K)_{\pi}$ is free over the localized Hecke algebra, which is a local complete intersection. We claim that it suffices to show that Hida’s Control Theorem, as stated in [36, Theorem 7.1], applies to $S_{\kappa(\pi)}(K)_{\pi}$. Indeed, the Hecke algebra ${\mathbb{T}}_{\bar{\rho}_{\nu},\emptyset}$ in weight $\kappa(\pi)$ is the quotient of ${\mathbb{T}}_{\pi}$ by the regular sequence corresponding to the weight $\kappa(\pi)$. Moreover, we are concerned with coherent cohomology in degree 00, so by Koecher’s principle the reduced localization hypothesis in Theorem 36 is superfluous.
Now it remains to verify that Hida’s Control Theorem does apply to weight $\kappa(\pi)$ when $\mu^{\prime}$ satisfies the regularity condition (ii) of Proposition 27. In fact, Boxer and Pilloni have proved a classicality theorem for overconvergent modular forms of small slope [8, Theorem 1.0.15 (3)] that asserts that overconvergent modular forms of weight $\mu$ on Shimura varieties of abelian type are classical if they satisfy a small slope condition and if $\mu^{\prime}$ satisfies the regularity condition (ii) of Proposition 27. Since ordinary modular forms automatically satisfy the small slope condition (see [8, Remark 1.0.6]), this completes the proof.